# Non-Expansive Mappings in Two-Time-Scale Stochastic Approximation: Finite-Time Analysis

### v1

@@ A.7 A Technical Lemma
Appendix B Proof for Theorem 2.2
Proof B.1 (Proof for Theorem 2.2).
@@
Lipschitz nature of $\hat{x}(\cdot)$
The map $\hat{x}(\cdot)$ is also Lipschitz with parameter $L_{0}=L/(1-\mu)$. This follows directly from the proof for Lemma 3.1, replacing $f$ with $f_{p}$.
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\left\|\mathcal{P}_{\mathcal{X}}\big{(}x_{k}+\alpha_{k}(f(x_{k},y_{k})-x_{k}+M_{k+1})\big{)}-\mathcal{P}_{\mathcal{X}}(\hat{x}(y_{k+1}))\right\|^{2}$
@@
This follows from the fact that $(x-f(x,y))$ is a $\mu^{\prime}$-strongly monotone function in $x$ for all $y$. This finally gives us
@@
The second term is dealt with as before. For the first term, we use (B) as follows
@@
Adding (B) and (B) as done in Lemma 3.3, we get that the iterates are bounded in expectation. For $\kappa_{1,p}=e^{D_{1}(\Gamma_{1,p}+\Gamma_{2})}$,
@@
Next, we prove almost sure convergence for our iterates. We have shown that the iterates satisfy that $\mathbb{E}[\|x_{k}\|^{2}]$ and $\mathbb{E}[\|y_{k}\|^{2}]$ are bounded by some constant and hence $\|x_{k}\|,\|y_{k}\|<\infty$ almost surely. We have also shown that $\hat{x}(y)$ is Lipschitz in $y$. For a fixed $y$, the ODE limit of the iteration is $\dot{x}(t)=f(x(t),y)-x(t)+\mathcal{B}_{\mathcal{X}}(x(t)).$ Here $\mathcal{B}_{\mathcal{X}}(x(t))$ is the boundary projection term required to keep $x(t)$ inside $\mathcal{X}$ at all times. Now, we also note that $\hat{x}(y)$ is the unique solution to the variational inequality VI$(\mathcal{X},x-f(x,y))$, given by the solution $x\in\mathcal{X}$ which satisfies $\langle z-x,x-f(x,y)\rangle\geq 0$ for all $z\in\mathcal{X}$. This follows from the fact that $x-f(x,y)$ is strongly monotone in $x$ for all $y$, and Proposition 1.5.8 and Theorem 2.3.3 from [15]. Then Theorem 7 from [13] states that the solutions of the ODE above converge to the solution of the VI, which is $\hat{x}(y)$ in this case. And hence using Lemma 8.1 from [2], $x_{t}\rightarrow\hat{x}(y_{t})$ with probability $1$. Now consider the ODE limit $\dot{y}(t)=g(\hat{x}(y(t)),y(t))-y(t).$ Theorem 12.2 and Remark 3 (mentioned below the theorem) from [2] together imply that $y_{t}$ almost surely converges to the set $\mathcal{\hat{Y}}$.
@@
Appendix C Proof for Theorem 4.1
Proof C.1 (Proof for Theorem 4.1).
@@
$\mathbb{E}\Big{[}x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\big{\rangle}\mid\mathcal{F}_{k}\Big{]}$
@@
Here $D_{6}=6L_{0}^{2}(1+1/\mu^{\prime})$ and $D_{7}=2\mathfrak{c}_{2}(1+3L_{0}^{2}+3L_{0}^{2}/\mu^{\prime})$. For appropriate choice of $C_{4}$ in our assumption that $\alpha_{0}\leq C_{4}$, we get $(\mu^{\prime 2}+6L_{0}^{2}L^{2}+6L_{0}^{2}L^{2}/\mu^{\prime})\alpha_{k}^{2}\leq\mu^{\prime}\alpha_{k}/2$.
@@
Here, $D_{8}=(L^{2}+1.5L_{J}L^{2})/\mu^{\prime}$. For appropriate choice of $C_{4}$ in our assumption that $\alpha_{0}\leq C_{4}$, we get $\beta_{k}^{2}\leq\beta_{k}\alpha_{k}$, and $(D_{6}D_{8}+1.5L_{J})\beta_{k}\alpha_{k}\leq 0.5\beta_{k}$.
@@
$\leq(\beta_{0}/\alpha_{0})\|x_{0}^{2}-x^{*}(y_{0})\|^{2}-(\beta_{k}/\alpha_{k})\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$
$\;\;-\sum_{i=1}^{k}\mathbb{E}[\|x_{i}-x^{*}(y_{i})\|^{2}]\left(\frac{\beta_{i-1}}{\alpha_{i-1}}-\frac{\beta_{i}}{\alpha_{i}}\right)\leq(\beta_{0}/\alpha_{0})\|x_{0}^{2}-x^{*}(y_{0})\|^{2}.$
Here, the final inequality follows from the fact that $\beta_{k}/\alpha_{k}$ is a decreasing sequence. Returning to (C), we have the assumption that the sequence $\alpha_{k}\beta_{k}$ is $\mathcal{O}(1/(k+1))$. We also have the assumption that the function $J(\cdot)$ is lower bounded. Hence the right-hand side of (C) is bounded by $D_{9}(1+\log(k+1))$ for some constant $D_{9}>0$. Hence,
@@
where $C_{3}=D_{9}/\beta_{0}$. The required $C_{4}$ in the assumption $\alpha_{0}\leq C_{4}$ for the result to hold is $C_{4}=\min\{1,(1-\mu)^{4}/(2+24L^{4}),(1-\mu)^{5}/(18(L^{6}+L^{5}+L^{4}+L^{2}+L))\}$.
Appendix D Proofs from Section 5: Applications
We first prove a lemma that relates strongly monotone and co-coercive operators with contractive and non-expansive mappings, respectively.
Lemma D.1.
@@ Appendix D Proofs from Section 5: Applications
D.1 Proof for Corollary 5.1
Proof D.3 (Proof for Corollary 5.1).
Since $H(x,y)$ is $\rho$-strongly-concave in $x$, $-\nabla_{x}H(x,y)$ is $\rho$-strongly monotone [22, Theorem 2.1.9]. Define $f(x,y)=x+(\rho/L^{2})\nabla_{x}H(x,y)$. Then $f(x,y)$ is $\mu-$contractive in $x$ for fixed $y$, where $\mu=\sqrt{1-\rho^{2}/L^{2}}$.
Recall that $\Phi(y)=\max_{x\in\mathbb{R}^{d_{1}}}H(x,y)$. Then Lemma 4.3 from [20] shows that $\nabla\Phi(y)=\nabla_{y}H(x^{*}(y),y)$, and $\nabla\Phi(y)$ is $2L_{0}L$-Lipschitz where $L_{0}=L/(1-\mu)$. Since $H(x,y)$ is convex in $y$, $\Phi(y)=\max_{x}H(x,y)$ is also convex in $y$. This implies that $\nabla\Phi(y)$ is $1/(2LL_{0})$ co-coercive [22, Theorem 2.1.5]. Now, let $g(x,y)=y-\frac{1}{LL_{0}}\nabla_{y}H(x,y)$, and hence $g(x^{*}(y),y)=y-\frac{1}{LL_{0}}\nabla\Phi(y)$. Hence $g(x^{*}(y),y)$ is non-expansive.
The iterates in (3) can then be written as
D.2 Proof for Corollary 5.2
Proof D.4 (Proof for Corollary 5.2).
Now, $\Delta=A_{22}-A_{21}A_{11}^{-1}A_{12}$ is non-zero PSD, and hence $\Delta$ is $1/\lambda_{\Delta,max}$-co-coercive [15, Pg. 79]. Then define $g(x,y)=y-(2/\lambda_{\Delta,max})(A_{21}x+A_{22}y-b_{2})$, and hence $g(x^{*}(y),y)$ is non-expansive. Here $x^{*}(y)=A_{11}^{-1}(-A_{12}y+b_{1})$.
Finally, the iterates in (5) can be rewritten as
D.3 Proof for Lemma 5.3
Proof D.5 (Proof for Lemma 5.3).
This implies that
$\leq L\|(AA^{T})^{-1}A\|\|x^{*}(\lambda_{1})-x^{*}(\lambda_{2})\|.$
Hence $\langle-Ax^{*}(\lambda_{1})+Ax^{*}(\lambda_{2}),\lambda_{1}-\lambda_{1}\rangle\geq\rho/(L\|(AA^{T})^{-1}A\|)^{2}\|\lambda_{1}-\lambda_{2}\|^{2}$. This implies that $-Ax^{*}(\lambda)$ is strongly monotone in $\lambda$.
b). Recall that $\hat{x}(\cdot)$ is the unique solution to $\mathcal{P}_{\mathcal{X}}(x+\nabla H_{0}(x)-A^{T}\lambda)=x$. Using [15, Prop. 1.4.2], $\hat{x}(\lambda)$ satisfies: $\left\langle x-\hat{x}(\lambda),\nabla H_{0}(\hat{x}(\lambda))-A^{T}\lambda\right\rangle\leq 0,\forall x\in\mathcal{X},\lambda\in\mathbb{R}^{d_{2}}$. Then we have
Combining these equations gives us
$-\left\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),\nabla H_{0}(\hat{x}(\lambda_{1}))-\nabla H_{0}(\hat{x}(\lambda_{2}))\right\rangle\leq-\left\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),A^{T}\lambda_{1}-A^{T}\lambda_{2}\right\rangle.$
Then, we have the following.
Now $\|A\hat{x}(\lambda_{1})-A\hat{x}(\lambda_{2})\|^{2}\leq\|A\|^{2}\|\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2})\|^{2}$, and hence $-A\hat{x}(\lambda)$ is $(\rho/\|A\|^{2})$-co-coercive.
D.4 Proof for Corollary 5.4
Proof D.6 (Proof for Corollary 5.4).
The iterates in (7) can be rewritten as

### v3

@@ A.7 A Technical Lemma
Appendix B Proofs from Section 5: Applications
B.1 Proof for Corollary 5.1
Proof B.1 (Proof for Corollary 5.1).
Since $H(x,y)$ is $\rho$-strongly-concave in $x$, $\nabla_{x}H(x,y)$ is $\rho$-strongly monotone [27, Theorem 2.1.9]. Define $f(x,y)=x+(\rho/L^{2})\nabla_{x}H(x,y)$. Then $f(x,y)$ is $\mu$-contractive in $x$ for fixed $y$, where $\mu=\sqrt{1-\rho^{2}/L^{2}}$.
Recall that $\Phi(y)=\max_{x\in\mathbb{R}^{d_{1}}}H(x,y)$. Then Lemma 4.3 from [24] shows that $\nabla\Phi(y)=\nabla_{y}H(x^{*}(y),y)$, and $\nabla\Phi(y)$ is $2L_{0}L$-Lipschitz where $L_{0}=L/(1-\mu)$. Since $H(x,y)$ is convex in $y$, $\Phi(y)=\max_{x}H(x,y)$ is also convex in $y$. This implies that $\nabla\Phi(y)$ is $1/(2LL_{0})$ co-coercive [27, Theorem 2.1.5]. Now, let $g(x,y)=y-\frac{1}{LL_{0}}\nabla_{y}H(x,y)$, and hence $g(x^{*}(y),y)=y-\frac{1}{LL_{0}}\nabla\Phi(y)$. Hence $g(x^{*}(y),y)$ is non-expansive.
The iterates in (3) can then be written as
B.2 Proof for Corollary 5.2
Proof B.2 (Proof for Corollary 5.2).
Now, $\Delta=A_{22}-A_{21}A_{11}^{-1}A_{12}$ is non-zero PSD, and hence $\Delta$ is $1/\lambda_{\Delta,max}$-co-coercive [19, Pg. 79]. Then define $g(x,y)=y-(2/\lambda_{\Delta,max})(A_{21}x+A_{22}y-b_{2})$, and hence $g(x^{*}(y),y)$ is non-expansive. Here $x^{*}(y)=A_{11}^{-1}(-A_{12}y+b_{1})$.
Finally, the iterates in (5) can be rewritten as
B.3 Proof for Lemma 5.3
Proof B.3 (Proof for Lemma 5.3).
This implies that
$\leq L\|(AA^{T})^{-1}A\|\|x^{*}(\lambda_{1})-x^{*}(\lambda_{2})\|.$
Hence $\langle-Ax^{*}(\lambda_{1})+Ax^{*}(\lambda_{2}),\lambda_{1}-\lambda_{1}\rangle\geq\rho/(L\|(AA^{T})^{-1}A\|)^{2}\|\lambda_{1}-\lambda_{2}\|^{2}$. This implies that $Ax^{*}(\lambda)$ is strongly monotone in $\lambda$.
[…]
Combining these equations gives us
$\left\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),\nabla H_{0}(\hat{x}(\lambda_{1}))-\nabla H_{0}(\hat{x}(\lambda_{2}))\right\rangle\leq-\left\langle\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2}),A^{T}\lambda_{1}-A^{T}\lambda_{2}\right\rangle.$
Then, we have the following.
Now $\|A\hat{x}(\lambda_{1})-A\hat{x}(\lambda_{2})\|^{2}\leq\|A\|^{2}\|\hat{x}(\lambda_{1})-\hat{x}(\lambda_{2})\|^{2}$, and hence $A\hat{x}(\lambda)$ is $(\rho/\|A\|^{2})$-co-coercive.
B.4 Proof for Corollary 5.4
Proof B.4 (Proof for Corollary 5.4).
The iterates in (7) can be rewritten as
Appendix C Proof for Theorem 2.2
Proof C.1 (Proof for Theorem 2.2).
@@ Appendix B Proofs from Section 5: Applications
[…]
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\left\|\mathcal{P}_{\mathcal{X}}\big(x_{k}+\alpha_{k}(f(x_{k},y_{k})-x_{k}+M_{k+1})\big)-\mathcal{P}_{\mathcal{X}}(\hat{x}(y_{k+1}))\right\|^{2}$
@@
This follows from the fact that (xf(x,y))(x-f(x,y)) is a $\mu^{\prime}$-strongly monotone function in $x$ for all $y$. This finally gives us
@@
The second term is dealt with as before. For the first term, we use (C) as follows
@@
Adding (C) and (C) as done in Lemma 3.3, we get that the iterates are bounded in expectation. For $\kappa_{1,p}=e^{D_{1}(\Gamma_{1,p}+\Gamma_{2})}$,
@@
[…]
@@
Appendix D Proof for Theorem 4.1
Proof D.1 (Proof for Theorem 4.1).
@@
$\mathbb{E}\Big[2\langle x_{k}-x^{*}(y_{k})+\alpha_{k}(f(x_{k},y_{k})-x_{k}),\alpha_{k}M_{k+1}+x^{*}(y_{k})-x^{*}(y_{k+1})\big\rangle\mid\mathcal{F}_{k}\Big]$
@@
Here $D_{6}=6L_{0}^{2}(1+1/\mu^{\prime})$ and $D_{7}=2\mathfrak{c}_{2}(1+3L_{0}^{2}+3L_{0}^{2}/\mu^{\prime})$. For appropriate choice of $C_{4}$ in our assumption that $\alpha\leq C_{4}$, we get $(\mu^{\prime 2}+6L_{0}^{2}L^{2}+6L_{0}^{2}L^{2}/\mu^{\prime})\alpha_{k}^{2}\leq\mu^{\prime}\alpha_{k}/2$.
@@
Here, $D_{8}=(L^{2}+1.5L_{J}L^{2})/\mu^{\prime}$. For appropriate choice of $C_{4}$ in our assumption that $\alpha\leq C_{4}$, we get $\beta_{k}^{2}\leq\beta_{k}\alpha_{k}$, and $(D_{6}D_{8}+1.5L_{J})\beta_{k}\alpha_{k}\leq 0.5\beta_{k}$.
@@
$\leq(\beta/\alpha)\|x_{0}^{2}-x^{*}(y_{0})\|^{2}-(\beta_{k}/\alpha_{k})\mathbb{E}\left[\|x_{k+1}-x^{*}(y_{k+1})\|^{2}\right]$
$\;\;-\sum_{i=1}^{k}\mathbb{E}[\|x_{i}-x^{*}(y_{i})\|^{2}]\left(\frac{\beta_{i-1}}{\alpha_{i-1}}-\frac{\beta_{i}}{\alpha_{i}}\right)\leq(\beta/\alpha)\|x_{0}^{2}-x^{*}(y_{0})\|^{2}.$
Here, the final inequality follows from the fact that $\beta_{k}/\alpha_{k}$ is a decreasing sequence. Returning to (D), we have the assumption that the sequence $\alpha_{k}\beta_{k}$ is $\mathcal{O}(1/(k+1))$. We also have the assumption that the function $J(\cdot)$ is lower bounded. Hence the right-hand side of (D) is bounded by $D_{9}(1+\log(k+1))$ for some constant $D_{9}>0$. Hence,
@@
where $C_{3}=D_{9}/\beta$. The required $C_{4}$ in the assumption $\alpha\leq C_{4}$ for the result to hold is $C_{4}=\min\{1,(1-\mu)^{4}/(2+24L^{4}),(1-\mu)^{5}/(18(L^{6}+L^{5}+L^{4}+L^{2}+L))\}$.
Appendix E Equivalence Between Formulations
We first show how a formulation using contractive and non-expansive maps implies the formulation using strongly monotone and co-coercive operators.
Lemma E.1.
Consider function $h:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$. Then,
  1. 1.
    If $h(\cdot)$ is $\mu$-contractive for $\mu\in[0,1)$, then function q(x)=xh(x)q(x)=x-h(x) is $(1-\mu)$-strongly monotone and $(1+\mu)$-Lipschitz.
  2. 2.
    If $h(\cdot)$ is non-expansive, then function q(x)=xh(x)q(x)=x-h(x) is $0.5$-co-coercive.
Proof E.2.
a). For $x_{1},x_{2}\in\mathbb{R}^{d}$,
$\langle x_{1}-x_{2},q(x_{1})-q(x_{2})\rangle=\|x_{1}-x_{2}\|^{2}-\langle x_{1}-x_{2},h(x_{1})-h(x_{2})\rangle.$
Now, $\langle x_{1}-x_{2},h(x_{1})-h(x_{2})\rangle\leq\|x_{1}-x_{2}\|\|h(x_{1})-h(x_{2})\|\leq\mu\|x_{1}-x_{2}\|^{2}.$ This implies that $\langle x_{1}-x_{2},q(x_{1})-q(x_{2})\rangle\geq(1-\mu)\|x_{1}-x_{2}\|^{2}$. The function $q(x)$ is trivially Lipschitz with constant $(1+\mu)$.
b). For $x_{1},x_{2}\in\mathbb{R}^{d}$, let s=xys=x-y and t=h(x)h(y)t=h(x)-h(y). Then, q(x)q(y)=stq(x)-q(y)=s-t. We first note that $2\langle s-t,s\rangle=\|s-t\|^{2}+\|s\|^{2}-\|t\|^{2}$. Since $h(\cdot)$ is non-expansive, $\|t\|\leq\|s\|$. This implies that $2\langle s-t,s\rangle\geq\|s-t\|^{2}$, and hence
$\langle q(x)-q(y),x-y\rangle\geq 0.5\|q(x)-q(y)\|^{2}.$
This completes the proof for Lemma E.1.
We now show the converse directions.
Lemma E.3.